Using identities, evaluate. 
71² ?????????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

71^2= (70+1)^2
= 70^2+ 140 + 1^2
= 4900 + 140 +1
= 5041

Thanks